Four coins are tossed.Let T be the random variable representing the number of tails that occur.find the values of the random variable T

1
Expert's answer
2022-02-22T12:42:22-0500

This is a binomial distribution with "n=4,p=\frac{1}{2}."

"P(T=0)=C_4^0(\frac{1}{2})^0(\frac{1}{2})^4=\frac{1}{16}."

"P(T=1)=C_4^1(\frac{1}{2})^1(\frac{1}{2})^3=\frac{1}{4}."

"P(T=2)=C_4^2(\frac{1}{2})^2(\frac{1}{2})^2=\frac{3}{8}."

"P(T=3)=C_4^3(\frac{1}{2})^0(\frac{1}{2})^1=\frac{1}{4}."

"P(T=4)=C_4^4(\frac{1}{2})^0(\frac{1}{2})^0=\frac{1}{16}."